suppose A=3i-2j and B=-j-aj ,what is the value of a such that A+B=2i


1
Expert's answer
2021-04-19T17:07:04-0400

Add each component of the vectors respectively:


"A+B=2i,\\\space\\
(3i-2j)+(-i-aj)=2i,\\
i-j(2+a)=2i."

The last is only possible when


"2+a=0, \\
a=-2."